The half-life of a certain radioactive substance is 15 years.
Suppose that at time t = 0, there are 29 g of the substance.
Then after t years, the number of grams of the substance remaining will be

How many grams of the substance will remain after 45 years?

Answer by ankor@dixie-net.com(15665)

The half life formula
A = Ao(2^(-t/h))
Where
A = resulting amt after t yrs
Ao = initial amt
h = half-life of substance
:
The half-life of a certain radioactive substance is 15 years.
Suppose that at time t = 0, there are 29 g of the substance.
Then after t years, the number of grams of the substance remaining will be:
A = 29(2^(-t/15))
:
How many grams of the substance will remain after 45 years?
A = 29(2^(-45/15))
A = 29(2^-3)
A = 29*.125
A = 3.625 grams